OpenStudy (summersnow8):

Need help finding correct answer (work included): A cyclist and her bicycle (combined mass of 75.0 kg) apply a force of 34.2 N. If the frictional force is -23.5 N, what will be the acceleration of the bicycle (in m/s2)? @rajat97

OpenStudy (summersnow8):

A cyclist and her bicycle (combined mass of 75.0 kg) apply a force of 34.2 N. If the frictional force is -23.5 N, what will be the acceleration of the bicycle (in m/s2)? not sure about this question m: 75.0 kg F = 34.2 N Ff: -23.5 N a: ? F = m * a 34.2 N = 75.0 kg * a a = 34.2 N Well, what is the net force on the bicycle, if cyclist plus bicycle exert force of 34.2 N and frictional force opposes to the tune of 23.5 N? From that net force, work out acceleration ΣF = 34.2 N - 23.5 N ΣF = 10.7 N F = m * a a = F / m a = 10.7 N / 75.0 kg a = 0.14266 m/s^2 a = 0.143 m/s^2

OpenStudy (nikato):

yup. that looks correct. i got the same thing.

OpenStudy (nikato):